# When Cars meet Drones: Hyperbolic Federated Learning for Source-Free Domain Adaptation in Adverse Weather

### v1

@@ 4.4 Server Aggregation (AG)
where $N_{k,c}^{r}=|\mathcal{F}_{k,c}|$ at round $r$ and $N_{c}^{r,tot}=\sum_{k\in\mathcal{K}}N_{k,c}$. We used the same smoothing parameter of Eq. 4, i.e., $\beta^{\prime}=\beta=0.85$.
Table 1: Data distribution ($\dagger$: with Cityscapes as clear, see Suppl. Mat. ) Clear Night Rain Fog SYNTH SELMA Train 24735 24735 24735 24735 Test 3087 3087 3087 3087 FLYAWARE-S Train 24771 24731 24808 24631 Test 3147 3139 3128 2935 REAL ${}^{\dagger}$ Train 1060 400 400 400 Test 214 100 100 100 FLYAWARE-R Train 94 35 35 35 Test 18 18 17 17 FLYAWARE-R-XL Train 565 565 565 565 Test 112 53 52 52
FA-S
FA-R
Refer to caption Day
Refer to caption Night
Refer to caption Rain
Refer to caption Fog
Figure 3: Samples from different datasets under varying weather conditions.
@@ 5 The FLYAWARE Adverse Dataset
In standard autonomous driving for semantic segmentation, there is a sufficient number of datasets providing a diversity of adverse conditions [42, 34]. However, in the case of aerial viewpoints, a scarcity of such datasets is evident, thus making the development of a dataset of this type a priority to unlock further research on the topic. To this aim, we introduce a new benchmark FLYAWARE - FLYing in Adverse Weather: Artificial and REal data. In the following section, we outline the structure of the synthetic and real aerial datasets, $\mathcal{D}_{drone}^{S}$ and $\mathcal{D}_{drone}^{T}$, respectively. They contain over 100K for which we contribute with 108180 generated images and 1287 domain-translated ones. In Tab. 1, we present the datasets used in this study. For additional information regarding the datasets for autonomous cars, please consult the Suppl. Mat.
FLYAWARE-S: Adverse Synthetic Dataset. Recently released, Syndrone [32] is a synthetic dataset based on the CARLA simulator [8]. While the dataset is richly annotated with 28 semantic classes, it currently lacks imagery in different weather conditions. We extended their work starting from the codebase provided in [32] by generating images in 3 adverse weather conditions (i.e., rain, fog, and night), while maintaining the capability of the system to produce images from multiple viewpoints and heights. The dataset includes drone views at heights ranging from 20 to 80 meters and with angles varying from 30 to 90 degrees, all with a resolution of 1920x1080. Moreover, we also generated 3D data (depth maps and LiDAR) for future usage in multimodal architectures.
FLYAWARE-R: Real Dataset Translation. Since no adverse weather dataset for aerial vehicles is available we opted for using image translation over standard drone datasets to build the adverse weather imagery. Addressing the task of converting daytime images into adverse conditions requires the availability of images of clear weather samples like sunny or cloudy conditions, and of adverse samples occurring in specific conditions of interest, such as rainy weather. Inspired by [13] which performs adverse domain translations for autonomous driving in the context of depth estimation, we opted for the use of Generative Adversarial Networks. For the rain and night adverse conditions, we employed the ForkGAN model [52] to translate clear-day training samples from the UAVid dataset [25]. While for the fog samples, we have used a combination of Omnidata [10] and FoHIS [49] methods. In the following, we will present an in-depth explanation of how we have performed the aforementioned conversion.
Day2Night. To convert daylight into nighttime, we trained ForkGAN over $14K$ images, half of which represent daytime and the other half nighttime. All the used images were sampled from Visdrone [53], and UAVDT [9], two datasets designed for Object Detection. We have decided to sample the training data from different datasets to increase the overall number of nighttime images. This was necessary to ensure that the GAN architecture reaches good enough reconstruction performance. The training phase lasted $40$ epochs. After the training, for all the nighttime samples in the UAVid dataset, we convert clear-day images into their nighttime counterparts using the pretrained model.
Day2Rain. Training the GAN models directly on drone data proved advantageous as it helped to mitigate domain shifts. In the rain case however there are no available drone datasets - not even for other tasks as in the case of night images -, therefore we had to train ForkGAN using diverse datasets that included adverse weather scenarios, like BDD100K [48], ACDC [34] and RainCityscapes [16] (which are datasets for automotive applications). For daytime to rainy conversion, we followed the same strategy applied for the conversion to nighttime training the ForkGAN model over $9K$ images for $40$ epochs and we applied the pretrained model to convert images into their rainy counterparts.
Day2Fog. To perform daytime to foggy conversion we first estimated a depth map for each image in the UAVid dataset using Omnidata [10], then, we exploited the estimated depth as input to the FoHIS method [49] to apply fog.
Federated Splits. We consider two scenarios: the first is an unbalanced setup with $40$ clients, corresponding to $32$ automotive clients and $8$ drones; the second is more balanced having $32$ automotive clients and $32$ drones, however the training and test set of the drones come from two different domains. Aligned with previous works [37], the clients have a different number of samples associated with them, in particular, the number of samples in each client is in the range $[69,72]$ for cars and $[24,25]$ for drones. To simulate real scenarios, we assumed that the most common weather would be clear daytime, while clients can have different combinations of adverse weather. Refer to the Suppl. Mat. for the weather distribution details on clients.
6 Implementation Details
Datasets. In this study, we leveraged two synthetic driving datasets - one collected from cars and one from drones - each encompassing a range of challenging illumination and weather conditions: the driving dataset SELMA [42] and the FLYAWARE-S dataset introduced in this work. Both datasets are based on the CARLA simulator [8] for image generation, making them ideal choices for our research due to their shared framework. Then, our method is adapted in a federated fashion to real-world datasets: for cars we selected the ACDC dataset, supplemented by Cityscapes [7] for the clear samples as described in the Suppl. Mat. As the drone counterpart, we used our adverse-weather translated dataset FLYAWARE-R, which offers labeled frames captured from aerial views. Additionally, to enhance the robustness of our approach, we performed also tests using the VisDrone object detection dataset for clients’ training as training at client side is unsupervised. In this setting the entire FLYAWARE-R dataset can be used for testing purposes (i.e. XL) for more reliable results. The adverse translation of FLYAWARE-R is described in Sec. 5. All the extended datasets and splits are referenced in Fig. 3 and will be available at the link: [soon]
Server pretraining. We chose DeepLabV3 architecture with MobileNetV2 as in [37]. We trained our model on the source synthetic dataset using a decreasing power-law learning rate $\eta$, starting at $\eta=5e-3$ with a power of 0.9. The optimization used SGD with momentum 0.9 and no weight decay, lasting for 5 epochs with batch size 16. For the weather classifier, we trained a 3-layer ConvNet with batch size 88, 8 epochs and SGD optimizer with learning rate $41e-$ 4.
Clients Adaptation. For the target dataset, experiments were run with a fixed learning rate of $\eta=1e-4$ with SGD optimizer. The training involved 5 clients per round for a total of $R=100$ rounds, with $\lambda_{cl}=140$. The pseudo-label teacher model was updated at the end of each round. For the optimizer of the manifolds, we use RiemmanianAdam as in [40] with $\gamma=0.1$ as initialization, learning rate equal to $31e-$ 3 and weight decay of $44e-$ 4.
Table 2: mIoU on the target datasets (ACDC and FLYAWARE-R (FA-R)).
@@ 6 Implementation Details
centralized FA-S+SELMA ACDC+FA-R Source Only 19.96 33.11 22.70
@@ 7 Experimental Results
The results are detailed in Tab. 2, which reports the average mIoU for cars, drones and a combined score. The latter is computed using the average for the classes present in both datasets and the car mIoU otherwise. For comparison, we initially examined 3 supervised settings: 1) the source only approach, which relies exclusively on synthetic labeled data from the source domains for training; 2) the fully centralized approach, where all available data is used to train the network specifically on the target domain ($250$ epochs, $lr=5e-3$); 3) the upper limit scenario for the federated model, where it undergoes fine-tuning with full supervision on the clients ($R=250$). In the unsupervised setting, (i.e., in the context of FFreeDA), we explore two strategies: the standard FedAvg [27] combined with our self-training approach, and the main competitor, LADD [37].
Table 7 and 7: Motivation of our approach: viewpoints generalization (Sec. 7); generalization to different weather conditions (Sec. 7).
[…]
Notably, unlike our strategy, LADD conducts dataset stylization from the target domain, assuming prior knowledge of the client data distribution during pretraining. We pretrain LADD on our synthetic adverse weather dataset to ensure comparability, maintaining the target’s average style as outlined in [37].
The large domain shift between the source and target domains is clear from the mIoU of $22.70\%$ achieved by the source only training compared to $58.08\%$ of the target training. Although the performance increases by almost $13\%$ by fine-tuning the source only pretraining, assuming labeled data on the client side is not very realistic. In the unsupervised scenario, the baseline approach of using self-teaching with pseudo-labels together with FedAvg leads to a score of $26.10\%$, with a relatively small improvement from the source only proving that simple FL strategies struggle in this challenging setting. LADD [37] allows to improve performances to $26.96\%$, an improvement of around $1\%$ compared to FedAvg. Our approach, in comparison to FedAvg and LADD, demonstrates a better improvement over the source only baseline reaching a mIoU of $28.96\%$. In particular notice how the gain is also well-distributed across car and drone agents, meaning that our strategy can effectively align with the intrinsic distribution shift that different viewpoints introduce. Notably, our approach is the only able to improve the source only results on both cars and drones, while competitors struggle with drone imagery since they are not able to generalize well to both and trade-off the better performances on cars with a reduction of the accuracy on drones. Qualitative results are reported in the Suppl. Mat.
Generalization capabilities of the proposed approach. We perform further tests to evaluate the capability of our approach to adapt to different viewpoints and weather conditions. Specifically, in Sec. 7 we can underline the following: 1) Jointly pretraining on both car and drones improves the overall mIoU of $4.47\%$ compared to car-only and $7.45\%$ to drone-only; 2) Adapting to car and drones together leads to roughly the same performances of adapting separately two models (one for cars and one for drones), i.e., the model generalize well across the two agents even when only one viewpoint is provided; 3) In the drone-only setting, the limited number of clients results in poor generalization and overconfidence in predictions. In Sec. 7, we underline that: 1) When pretraining with synthetic data from all weather conditions the model improves on all, especially the night; moreover, adapting from clear weather only results in poor generalization when adapting to real data; 2) Although adapting with only clear weather data increases the mIoU on clear weather itself, it worsens the performance on almost all the other conditions.
Table 3: Ablation of the proposed method components. $\mathcal{L}_{st}$ AG $\mathcal{L}_{cl}$ BN Car Drone All 19.96 33.11 22.70 25.00 30.25 26.10 24.75 33.49 26.57 24.91 34.55 26.92 27.17 33.60 28.51 27.45 34.69 28.96
Figure 4: Training curves of the ablation.
@@
Ablation studies. Tab. 3 shows an extensive ablation of the various modules of the proposed approach together and Fig. 4 the corresponding training curves. The aggregation strategy of Eq. (12) stabilizes the self-training and avoids performance decrease on unbalanced clients (i.e. drones). The batch normalization effect is visible in cars and drones ($+3.4\%$ overall). Observe that the testing set of drones is more balanced in terms of weather categories (¼  each), making it more noticeable. The hyperbolic prototypes allow for improving the generalization of the feature spaces compared to the baseline, achieving $28.51\%$. In Suppl. Mat. a comprehensive examination of the advantages over the standard Euclidean space is provided. Finally, adding each component leads to the final score of $28.96\%$. Overall, the contributions of all the components jointly produce a gain of over $6.26\%$ of mIoU on the source only. Further ablation on the hyperparameters is in the Suppl. Mat.
@@ 8 Conclusion
In this paper we introduced a new federated learning scenario, addressing the challenges of heterogeneous autonomous agents in adverse weather conditions. The contributions include a realistic federated setting where clients are unsupervised and a versatile model seamlessly integrating aerial views. To achieve these results, we introduced the FLYAWARE dataset for semantic segmentation in adverse weather. The proposed method, HyperFLAW, employs weather-aware batch normalization layers to mitigate domain shifts across diverse weather conditions. Additionally, it leverages prototype-based learning in the hyperbolic space, contributing a novel method for consistent training in this challenging federated setting. By extending the scope to different types of vehicles and addressing real-world challenges such as adverse weather and privacy preservation, our contributions lay a foundation for future advancements in autonomous driving. Further research will be devoted to the exploitation of multi-modal data and to more advanced ad-hoc adaptation strategies for adverse weather conditions.
Acknowledgment
@@
This work was supported in part by the European Union through the Italian National Recovery and Resilience Plan (NRRP) of NextGenerationEU, a partnership on “Telecommunications of the Future” (Program “RESTART”) under Grant PE0000001.
Supplementary Material
Giulia Rizzoli* Matteo Caligiuri* Donald Shenaj Francesco Barbato Pietro Zanuttigh
**footnotetext: These authors contributed equally to this work.
This document contains the supplementary material for the paper When Cars meet Drones: Hyperbolic Federated Learning for Source-Free Domain Adaptation in Adverse Weather. We start by introducing additional details on the employed datasets. We then proceed with ablations supporting the usage of the weather batch-normalization and of hyperbolic prototypes. Finally, we show further experimental data including per-class accuracy scores and qualitative results.
S1 Data Selection and Distribution
In the main paper, we introduced the FLYAWARE aerial dataset since it is a novel contribution. Conversely, we offered fewer details for car driving datasets, as we relied on existing datasets. Here we will first describe the technical choices adopted to enable distributed learning on car agents in adverse weather. Then, some additional details on weather distribution among clients are provided.
S1.1 Driving Datasets
For the synthetic source dataset for cars, we used the SELMA dataset [42]. It offers a comprehensive set of 27 weather and daytime conditions, resulting in a vast dataset of over 20M samples. To better align with the weather scenarios considered in real data [34], we opted not to use the standard SELMA split, but downloaded over 24k samples in the ClearNoon, ClearNight, HardRainNoon, MidFoggyNoon splits. In total, for this work, we employed almost 100k SELMA samples from the Desk Cam point of view that match the one used in the real-world dataset. As the real counterpart, we used the ACDC dataset, from which we selected the 3 domains — night, rain, and fog — that match our pretraining. As ACDC lacks images in clear weather, we supplemented it with daytime images coming from the Cityscapes dataset [7], which is the most similar to ACDC. Note that, the ACDC dataset has been built to create an adverse condition version of Cityscapes, and shares the same class-set. Cityscapes provides more samples than those in the thematic splits by ACDC. Therefore, we subsampled its training set to match the sizes. Since the goal was to select clear weather conditions, we devised an automatic way of assigning a sunlit level (Eq. S1) to each image and selected the needed images by sorting them according to this metric:
$\text{sunlit}(\textbf{X})=\sum\limits_{c\in\{r,g,b\}}\frac{1}{HW}\sum\limits_{i=0}^{W}\sum\limits_{i=0}^{W}\textbf{X}[i,j,c]$ (S1)
As shown in Tab. 1 in the main paper, we selected a set of images to match the total count of all other conditions combined, aiming for a ratio of three clear sky images to one adverse condition. This decision was driven by the desire to retain “clear sky” as the most probable scenario in real-world contexts. In Fig. S1 we report some examples of samples with their rating (a bright one, a dark one and a mid-range one).
Figure S1: Cityscapes samples and corresponding sunlit score. Decreasing from the left, high/mid/low-score ($\sim 300/200/100$).
S1.2 Weather Heterogeneity
Fig. S2 illustrates the weather distribution among clients in the configuration ACDC+ FLYAWARE-R. We deliberately supplied clear day samples to each client to emulate a common scenario where clear daylight conditions prevail, with fewer instances of challenging weather conditions. This introduces an inherent challenge as adverse weather data samples are less prevalent. Additionally, the imbalance among clients further complicates the scenario.
Figure S2: Distribution of weather conditions across clients (ACDC and FLYAWARE-R).
Fig. S3 shows the weather distribution for clients in the ACDC+FLYAWARE-R-XL setup. While this configuration exhibits better data balance across clients, it introduces new complexities: some drone clients exclusively operate in adverse conditions, and the training and testing domains originate from different datasets (VisDrone [53] and UAVid [25], respectively). Notably, VisDrone images offer real night data, whereas we utilized domain-translated images for fog and rain using the same strategies described in the main paper.
Figure S3: Distribution of weather conditions across clients (ACDC and FLYAWARE-R-XL).
S2 Impact of Weather-Batch Normalization
As shown in Tab. 3 of the main paper, the inclusion of ad-hoc batch norms (BNs) enhances performance in comparison to utilizing a non-personalized network. First of all, in Fig. S4, we show in detail the architecture of the weather classifier. Although the classifier model is simple, it achieves an accuracy of $98.96\%$ on the source synthetic datasets used to train it and of $72.37\%$ on the real world target ones.
Figure S4: Weather classifier architecture model.
Table S1: Effect of the different modules on the weather.
Modules Clear Night Rain Fog
Pretrain 25.41 13.62 26.72 26.08
ST + AG w/o BN 29.27 13.47 29.47 27.86
ST + AG w BN 26.81 13.50 30.19 28.86
ALL w/o BN 31.58 14.04 30.53 28.93
ALL w BN 31.73 14.18 34.19 29.86
We further examine in Tab. S1 the advantage of adapting the system to accommodate different weather conditions as a remarkable feature for the decision-making of autonomous driving agents. The basic self-training strategy, coupled with the proposed server-side aggregation scheme, has shown stability and an overall improvement over pretraining. However, due to the prevalence of clear-day images in most clients’ samples, the network tends to learn more about this weather condition. Introducing the personalized weather BNs helps to mitigate this, improving performances in adverse weather conditions, with an mIoU increase of $3\%$ on Rain and $1.78\%$ on Fog. Similar results are evident when comparing the full method with and without the weather BNs, showcasing mIoU increases of $3.66\%$ on Rain and $0.93\%$ on Fog. As a side note, the most difficult scenario remains the night where the improvement over the pretraining is just $0.56\%$ of mIoU. Overall, we remark on the significance of incorporating personalized weather BNs for mitigating bias towards clear day images and enhancing performance across various weather conditions.
As a final note, considering that modern vehicles often have access to real-time weather information through several sensors (e.g., automatic windshield wipers for rain detection, clock or optical light detection for night and fog), and the weather information might be directly obtained.
S3 Analysis of class prototypes effectiveness
Hyperbolic vs. Euclidean Prototypes
Figure S5: Euclidean vs hyperbolic prototypes.
In Fig. S5, we show the training curves for both Euclidean prototypes and hyperbolic ones employing different values of the weighting parameter for the prototype loss $\lambda_{cl}$. Notably, hyperbolic space prototypes consistently allow for an increase in performances across different $\lambda_{cl}$ values. The key strength of the hyperbolic space lies in its ability to effectively represent distances between feature vectors, a crucial factor in tasks like clustering where having a distance function capturing meaningful relationships among data points is essential.
Setting of the $\lambda_{cl}$ parameter Fig. S6 shows the mIoU for different values of the $\lambda_{cl}$ parameter. Best performances are achieved in the range 140-180, with a quite stable maxima.
Figure S6: Tuning on $\lambda_{cl}$ weight.
S4 Per-class Results
Tab. S2 contains a more detailed version of the results in Tab. 3 of the main paper that shows also the class-by-class accuracy. As common on the employed datasets, the mIoU is higher in the common classes (e.g., road or building) and lower in the rare and more challenging ones. However, it can be seen that by adding the various components of the model, results tend to increase consistently in most classes, even if a few challenging ones remain hard to detect. Some classes like traffic sign, car or bicycle show impressive improvements. On the other side, there are a few cases in which our approach does not improve the results on some under-represented classes such as terrain and rider.
Table S2: Per class IoU for the proposed approach and its ablated versions.
ST
AG
CL
BN
road
sidewalk
building
wall
fence
pole
traffic light
traffic sign
vegetation
terrain
sky
person
rider
car
truck
bus
train
motorcycle
bicycle
All
35.4 5.7 55.8 4.6 4.2 14.2 13.3 10.9 61.9 6.3 43.7 26.0 17.2 11.1 3.3 3.8 1.4 8.9 10.7 22.7
43.7 22.0 57.2 6.6 4.7 15.8 9.8 13.9 50.2 5.7 68.4 24.3 11.4 34.6 8.9 4.7 3.5 16.0 14.9 26.1
43.9 23.2 59.2 7.1 5.9 16.2 11.1 14.8 55.3 6.1 58.5 26.0 11.9 34.2 9.2 4.3 3.0 15.8 13.5 26.6
43.8 23.6 59.7 6.9 5.9 15.9 10.7 14.8 56.5 5.1 60.6 25.4 11.9 37.7 9.0 4.5 2.8 15.1 13.0 26.9
46.8 22.4 59.9 7.4 6.7 17.6 17.0 22.4 56.6 7.1 62.6 25.4 10.5 35.5 9.5 6.9 3.7 16.6 25.7 28.5
47.2 22.7 60.6 7.0 6.7 17.3 16.8 22.7 57.7 5.9 64.9 25.4 10.6 36.9 9.3 6.8 3.5 16.8 25.5 29.0
S5 Qualitative Results
Fig. S7 shows some qualitative results for car and drone samples in both clear and adverse weather. FedAvg [27] and LADD [37] appear to be overconfident on the road and car classes, respectively. While our model adeptly captures the structure of the street better than other models in both adverse and non-adverse conditions, the adverse conditions images remain a challenging aspect. For drones, the discrepancies between the predicted classes and the ground truth arise also due to the fact that the network is trained on a superset of classes beyond those present in the drone dataset. Notice that sometimes the network assigns “fine” classes that share a semantic affiliation with the respective “coarse” classes, e.g., predict terrain instead of vegetation, but on the drone dataset labeling terrain is not present and terrain samples have ground truth set to vegetation. The same thing happens for the sidewalk and road or rider and person. For this reason, we remapped the 19 classes of the full set into the 5 drone ones using the mapping proposed in [32]. The figure shows both the original prediction maps from the network and the ones obtained after remapping the 19 classes into the 5 of the drone datasets. The comparison with the ground truth shows that some predictions not matching are just due to the different types of labeling in drone and car datasets.
RGB GT FedAvg[27]+ST LADD[37] HyperFLAW
CAR CLEAR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DRONE CLEAR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DRONE CLEAR (remapped) Refer to caption Refer to caption Refer to caption
CAR ADVERSE Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DRONE ADVERSE Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DRONE ADVERSE (remapped) Refer to caption Refer to caption Refer to caption
Road Sidewalk Building Wall Fence Pole T. Light T. Sign Vegetation Terrain
Sky Person Rider Car Truck Bus Train Motorbike Bicycle Unlabeled
Figure S7: Qualitative results.

### v2

@@ 4.4 Server Aggregation (AG)
where $N_{k,c}^{r}=|\mathcal{F}_{k,c}|$ at round $r$ and $N_{c}^{r,tot}=\sum_{k\in\mathcal{K}}N_{k,c}$. The smoothing parameter is the same of (4), $\beta^{\prime}=\beta=0.85$.
Clear Night Rain Fog
SYNTH SELMA Train 24735 24735 24735 24735
Test 3087 3087 3087 3087
FLYAWARE-S Train 24771 24731 24808 24631
Test 3147 3139 3128 2935
REAL ACDC Train 1060 400 400 400
Test 214 100 100 100
FLYAWARE-R Train 94 35 35 35
Test 18 18 17 17
FLYAWARE-R-XL Train 565 565 565 565
Test 112 53 52 52
Table 1: Data distribution. $\dagger$: Cityscapes as clear, see Suppl. Mat.
FLYAWARE-S
FLYAWARE-R
Day
Night
Rain
Fog
Figure 5: FLYAWARE: samples under different weathers.
@@ 5 The FLYAWARE Adverse Dataset
In standard autonomous driving for semantic segmentation, there is a sufficient number of datasets providing a diversity of adverse conditions [37, 31]. However, in the case of aerial viewpoints, a scarcity of such datasets is evident, thus making the development of a dataset of this type a priority to unlock further research on the topic. To this aim, we introduce a new benchmark FLYAWARE - FLYing in Adverse Weather: Artificial and REal data. In the following section, we outline the structure of the synthetic and real aerial datasets, $\mathcal{D}_{drone}^{S}$ and $\mathcal{D}_{drone}^{T}$, respectively. They contain over 100K images for which we contribute with 108180 generated images and 1287 domain-translated ones. Fig. 5 shows samples from FLYAWARE. Please consult the Suppl. Mat. for details on the dataset construction.
@@
centralized SELMA+FA-S ACDC+FA-R Source Only 19.96 33.11 22.70
@@
(a) mIoU on the target datasets (ACDC and FLYAWARE-R (FA-R))
Pretrain on Source Only Adapted
Separately Jointly
Car+Drone 22.70 28.95 28.96
Car only 18.23 25.17 25.24
Drone only 15.25 14.37 13.68
(b) Viewpoint Generalization
Pretrain Adapt Clear Night Rain Fog All
Clear - 23.43 6.43 24.93 25.50 18.28
All - 25.41 13.62 26.72 26.08 22.69
Clear Clear 26.87 1.02 8.02 9.86 14.36
Clear All 25.97 7.87 25.70 23.16 18.11
All Clear 30.96 10.85 22.22 26.38 25.03
All All 31.73 14.18 34.19 29.86 28.96
(c) Weather conditions’ effects
Table 2: Main results of the work over the multi-view setting. Tab. 2 reports the mIoU on multiple target datasets. Tab. 2 shows a viewpoint generalization study where during adaptation, clients train either two separate models or jointly train a single model. Tab. 2 presents a study on the generalization to different weather conditions.
Datasets. In Tab. 1, we present the datasets used in this study. Our research leverages two synthetic driving datasets based on the CARLA simulator [10]: the car-centric SELMA dataset [37] and our newly introduced drone-centric FLYAWARE-S dataset. Both encompass a range of challenging illumination and weather conditions. Additionally, although lacking aerial counterpart, we include GTAV [29] dataset in the Suppl. Mat. for comparison with previous supervised federated methods.
We then adapt our method to real-world datasets from different domains in a federated fashion. For cars, we utilize the ACDC [31] dataset, supplemented by Cityscapes [8] for clear weather samples (detailed in Suppl. Mat. ). For drones, we created FLYAWARE-R, an adverse-weather translated dataset, using two source datasets: UAVid [23], which is labeled for segmentation but limited in size, and VisDrone [47], which is intended for object detection. In FLYAWARE-R, we train and test on UAVid domain-translated samples. For enhanced robustness, we also created an extended set (XL) where we train on VisDrone domain-translated samples and evaluate on UAVid, allowing for more comprehensive testing. The adverse weather translation process is described in Suppl. Mat. All datasets and splits are referenced in Tab. 1.
Federated Scenarios. We examine two main federated learning settings: (i) an Unbalanced setting with 40 clients (32 automotive, 8 drones); (ii) a Balanced one with 64 clients (32 automotive, 32 drones), with drone training and test sets from different domains. Following [34], we sampled a heterogeneous distribution across clients: automotive clients have 69-72 samples each, while drone clients have 24-25 samples. To simulate real-world conditions, we set clear daytime as the predominant weather, with clients experiencing various combinations of adverse conditions. For weather distribution across clients, see the Suppl. Mat. Additionally, we explored a limit-case scenario: (iii) 64 balanced clients (as in (ii)), each assigned a single weather condition, to isolate weather variability effects.
Table 3: Ablation studies. Tab. 3 compares different BN strategies in three different scenarios: (Scenario (i)) where clients are in ratio 4:1 (car:drones); (Scenario (ii)) with ratio 1:1; (Scenario (iii)) where clients (in 1:1 ratio) have data from a single weather condition. Tab. 3 reports a more complete ablation study on all components of our strategy.
Method Aggr. W-BNs (i) (ii) (iii)
SiloBN [1] 25.09 24.50 24.93
FedBN [20] 25.20 24.70 25.44
Ours 28.51 27.07 25.49
Ours 28.96 28.84 28.04
(a) Batch normalization strategies.
$\mathcal{L}_{st}$ AG $\mathcal{L}_{cl}$ BN Car Drone All
19.96 33.11 22.70
25.00 30.25 26.10
24.75 33.49 26.57
24.91 34.55 26.92
27.17 33.60 28.51
27.45 34.69 28.96
(b) Ablation on components.
Figure 6: Euclidean (dashed lines) vs hyperbolic (solid lines) prototypes.
Figure 7: Training curves of the ablated method.
@@ 6 Experimental Results
The results are detailed in Tab. 2, which reports the average mIoU for cars, drones and a combined score. The latter is computed using the average for the classes present in both datasets and the car mIoU otherwise. See the Suppl. Mat. for implementation and network architecture details. For comparison, we initially examined three supervised baselines: 1) the source-only approach, which relies exclusively on synthetic labeled data from the source domains for training; 2) the fully centralized approach, where all available data is used to train the network specifically on the target domain; 3) the upper limit scenario for the federated model, where it undergoes fine-tuning with full supervision on the clients. In the unsupervised setting, (i.e., in the context of FFreeDA), we explore two strategies: the standard FedAvg [25] combined with our self-training approach, and the main competitor, LADD [34]. Notably, unlike our strategy, LADD conducts dataset stylization from the target domain, assuming prior knowledge of the client data distribution during pretraining. We pretrain LADD on our synthetic adverse weather dataset to ensure comparability, maintaining the target’s average style as outlined in [34]. The significant domain shift between source and target domains can be observed from the mIoU of 22.70% achieved by source-only training compared to 58.08% of the target training. Although fine-tuning the source-only pretraining increases performance by almost 13%, assuming labeled data on the client side is unrealistic in many scenarios. Our approach can successfully align the intrinsic distribution shift that different viewpoints introduce, in addition to outperforming other methods. The Suppl. Mat. demonstrates this result holds even when no aerial data are provided in the GTA $\rightarrow$ Cityscapes scenario.
Multi-viewpoint generalization capabilities. Tab. 2 shows that multi-viewpoint pretraining enhances overall performance compared to pretraining on either car or drone data alone. Specifically, jointly pretraining on both cars and drones improves the overall mIoU by 4.47% compared to car-only and 7.45% to drone-only. Importantly, we demonstrate that clients of the other agent type are not necessary for training. Adapting to cars and drones together leads to roughly the same performance as adapting two separate models - one for cars and one for drones -, indicating good generalization across the two agents even when only trained with data from a single domain.
Impact of Weather Conditions and Normalization. In Tab. 2, we present experiments training on the clear-only weather subset to verify generalization abilities across weather conditions. When performing supervised pretraining and unsupervised adaptation (i.e., our full method) on clear-only data and then testing on adverse conditions, the model loses over 14 mIoU points, motivating further exploration of this setting. To understand the impact of batch-normalization - our main component to address climate heterogeneity - we also tested a limit-case scenario where each client has data of only one weather condition (Tab. 3, Scenario (iii)). In this context, the effect of our weather-specific BNs is enhanced, demonstrating that sharing weather-personalized features benefits in learning a global model. Note that not aggregating BN layers, as in SiloBN [1] and FedBN [20], worsens the performance over the baseline (i.e., aggregating over non-weather specific BNs), proving that these layers carry weather-specific style content. Further per-weather results are discussed in the Suppl. Mat.
Impact of prototypes in the Hyperbolic space. The key strength of the hyperbolic space lies in its ability to effectively represent distances between feature vectors, a crucial factor in tasks like clustering where having a distance function capturing meaningful relationships among data points is essential. In Fig. 6, we show the training curves for both Euclidean and Hyperbolic prototypes employing different values of the weighting parameter for the prototype loss $\lambda_{cl}$. Hyperbolic space prototypes consistently allow for increased performance across different $\lambda_{cl}$ values.
@@
Ablation studies. Tab. 3 shows an extensive ablation of the various modules of the proposed approach together and Fig. 7 the corresponding training curves. The aggregation strategy of Eq. (12) stabilizes the self-training and avoids performance decrease on unbalanced clients (i.e. drones). The batch normalization effect is visible in cars and drones ($+3.4\%$ overall). Observe that the testing set of drones is more balanced in terms of weather categories (¼  each), making it more noticeable. The hyperbolic prototypes allow for improving the generalization of the feature spaces compared to the baseline, achieving $28.51\%$. Finally, adding each component leads to the final score of $28.96\%$. Overall, the contributions of all the components jointly produce a gain of over $6.26\%$ of mIoU on the source only. Further ablation on the hyperparameters is in the Suppl. Mat.
@@ 7 Conclusion
In this paper we introduced a new federated learning scenario, addressing the challenges of heterogeneous autonomous agents in adverse weather conditions. The contributions include a realistic federated setting where clients are unsupervised and a versatile model seamlessly integrating aerial views. We also introduced the FLYAWARE dataset for semantic segmentation in adverse weather. The proposed method, HyperFLAW, employs weather-aware batch normalization layers to mitigate domain shifts across diverse weather conditions. Additionally, it leverages prototype-based learning in the hyperbolic space, contributing a novel method for consistent training in this challenging federated setting. By extending the scope to different types of vehicles and addressing real-world challenges such as adverse weather and privacy preservation, our contributions lay a foundation for future advancements in autonomous driving. Further research will address the exploitation of multi-modal data and explore more advanced adaptation strategies for adverse weather conditions, as well as bridging the significant domain gap between car and aerial images.
@@
Acknowledgment.  This work was partially supported by the European Union under the Italian National Recovery and Resilience Plan (NRRP) of NextGenerationEU, partnership on “Telecommunications of the Future” (PE00000001- program “RESTART”).